OpenStudy (anonymous):

I hate this part of squareroots!!! please help I don't get it 4√12-7√27

OpenStudy (anonymous):

learn to be okay with the properties of exponents/roots \[\sqrt{ab}=\sqrt{a} \cdot \sqrt{b}\]

OpenStudy (anonymous):

you are multiplying 4 times the square root of 12 first. what is it? use calculator.

OpenStudy (anonymous):

look for factors that are perfect squares... \[12 = 3\cdot 4\] \[27 = 3\cdot 9\]

OpenStudy (anonymous):

@peoplesay1214 why use a calculator? it's not needed. if so, only after simplifying... and this one does simplify.

OpenStudy (anonymous):

I know its easy to do it on calculator but I just want to make sure I know how to do it in my head for tests/quizes lol but here is how I attempted it.... 4√3 - 3√9 = -9+4√3=-13√3?

OpenStudy (anonymous):

oh yes simplify... then solve

OpenStudy (anonymous):

almost... like this: \[4\sqrt{12}-7\sqrt{27}=4\sqrt{4}\cdot \sqrt{3}-7\sqrt{9}\cdot \sqrt{3}=4\cdot2 \sqrt{3} - 7\cdot 3 \sqrt{3}\] \[8\sqrt{3}-21\sqrt{3}=\left( 8-21 \right)\sqrt{3} = -13 \sqrt{3}\] remember, it's about getting the process right... then you're learning. if you only get the answer right. you're only getting lucky if you're not getting the process.

OpenStudy (anonymous):

Oh thank you so much I so get it now. I have a crappy professor and this part what like foreign language lol

OpenStudy (anonymous):

math often is!!!